OpenStudy (anonymous):

how do I express this expression as a single logarithm: (1/2)[log3(a) + 2log5(b)] - 3log5(c)]?

OpenStudy (anonymous):

here's what I have: = (1/2)[log3(a) log5(b)^2] - log5(c)^3] = 1/2log3(a) + 1/2log5(b)^3 - log5(c)^3 = log3(a)^1/2 + log5(b)^3x1/2 - log5(c)^3 = log3(a)^1/2 + log5((b)^3/2 / c^3)

OpenStudy (anonymous):

OK....you need first collect all terms with the same base 2log5(b)-3log5(c)=log5(b)^2-log5(c)^3=log5[(b^2)/(c^3)]

OpenStudy (anonymous):

oh never mind! I got it, I mistakenly put log3 instead of log5!

OpenStudy (anonymous):

oh ok